MASSACHUSETTS INSTITUTE OF TECHNOLOGY ARTIFICIAL INTELLIGENCE LABORATORY

A.I. Memo No. 848b November 1991

Revised^{4} Report on the Algorithmic Language Scheme

William Clinger and Jonathan Rees (Editors)

H. Abelson  R. K. Dybvig  C. T. Haynes  G. J. Rozas
N. I. Adams iv D. P. Friedman  E. Kohlbecker  G. L. Steele Jr.
D. H. Bartley R. Halstead  D. Oxley  G. J. Sussman
G. Brooks  C. Hanson  K. M. Pitman  M. Wand

Abstract

Data and procedures and the values they amass, Higher-order functions to combine and mix and match, Objects with their local state, the messages they pass, A property, a package, the control point for a catch— In the Lambda Order they are all first-class. One Thing to name them all, One Thing to define them, One Thing to place them in environments and bind them, In the Lambda Order they are all first-class.

Keywords: Scheme, Lisp, functional programming, computer languages.

This report describes research done at the Artificial Intelligence Laboratory of the Massachusetts Institute of Technology. Support for the laboratory’s artificial intelligence research is provided in part by the Advanced Research Projects Agency of the Department of Defense under Office of Naval Research contract N00014-89-J-3202 and by the National Science Foundation under grant number MIP-9001651.

With apologies to J. R. R. Tolkien.